At what temperature (in degrees Celsius) will xenon atoms have the same average speed that Cl2 molecules have at 20 ∘C?
Answer #1

1/2Mv2 = 3/2RT


from this
v2 = 3/2RT x 2 / M
v2 = 3RT/M
since the speeds are going to be the same...
3RT1/M1 = 3RT2/M2
cancel out 3R and get ...
T1/M1 = T2/M2
solve for T2
T2 = T1M2/M1
T2 = 293K x 131.3g/mol / 70.9 g/mol
T2 = 542.6K

or 542.6 - 273

= 270 ºC